# Information Elicitation in Agency Games

### v2

@@ 3.1.2 Revealed information contract
For ease of exposition, our analysis will focus on a one-dimensional binary feature: $\mathcal{X}=\{0,1\}$, where $X=1$ with probability $\theta$. We frame the principal’s decision problem as that of choosing $\rho(0)=p_{0}$ and $\rho(1)=p_{1}$. Further extensions to higher cardinalities or continuous $X$ may be of practical interest, though we expect our qualitative insights to extend to these settings.
@@
The principal moves first and chooses $p_{0}^{*},p_{1}^{*}\in\arg\max_{p_{0},p_{1}}\Pi_{\text{rev}}(p_{0},p_{1})$, resulting in an equilibrium where the agent’s utility is $V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$.
@@
In the rest of this section, we will analyze the effects of information revelation on the agent’s utility, principal’s utility, and total welfare. Specifically, we will compare $V_{\text{con}}(p^{*})$ to $V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$ to understand the agent’s incentive to reveal $X$. We will also analyze the consequences of the agent’s revelation decision on the principal’s utility, as well as consequences on total welfare.
3.2 Agent’s Revelation Incentives
The central goal in this work is to analyze the circumstances under which the agent would prefer to either conceal or reveal the existence of $X$ at time $t=1$. Thus, to begin, we will analyze situations that lead to $V_{\text{con}}(p^{*})$ being higher than $V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$ or vice versa, focusing on properties of the distributions $F$, $F_{0}$, and $F_{1}$. In the analogy to price discrimination, we may think of $F(p)$ as the task completion “quantity” as a function of price $p$, or the proportion of agents drawn uniformly at random from a population with costs distributed as $C$ that would complete the task for price $p$.
We present sufficient conditions on the distributions $F_{0},F_{1}$ for the agent to either prefer to conceal or to prefer to reveal. We then build intuition for these conditions through an example using exponential and Weibull distributions.
3.2.1 Concealment condition with one zero-cost type
@@
First, we present a sufficient condition for the agent to prefer to conceal $X$ when one of the agent types is anchored at zero. That is, $X=1$ implies that the agent incurs zero cost. Proposition 1 gives a sufficient condition on $F_{0}$ for the agent to prefer for the environmental variable $X$ to remain concealed.
@@
Qualitatively, the elasticity $\eta$ captures the sensitivity of the task completion to price. Thus, the inequality in equation (3) corresponds to a scenario when the sensitivity of the task completion to price when $X=0$ does not differ too strongly from that of the concealed setting. For example, this arises when $F_{0}$ is close to the constant function 1. We give another example using the exponential distribution in Section 3.3 below, where equation (3) holds if the mean of $F_{0}$ is low enough. In summary, the agent prefers concealment if the higher cost type still has relatively low cost.
3.2.2 Concealment and revelation conditions under a decreasing ratio assumption
To give additional sufficient conditions for concealment and revelation beyond the anchored setting with one zero-cost type, we apply an analysis technique similar to that of Aguirre et al. [2010], who analyzed the effects of third degree monopoly price discrimination on total welfare.

### v3

@@ Appendix A Additional Notation
Some of our analysis will focus on a one-dimensional binary feature: $\mathcal{X}=\{0,1\}$, where $X=1$ with probability $\theta$. In these cases, we simplify notation by parameterizing the principal’s decision problem as that of choosing $\rho(0)=p_{0}$ and $\rho(1)=p_{1}$.
@@
The principal moves first and chooses
$p_{0}^{*},p_{1}^{*}\in\arg\max_{p_{0},p_{1}}\Pi_{\text{rev}}(p_{0},p_{1}),$
resulting in an equilibrium where the agent’s utility is $V_{\text{rev}}(p_{0}^{*},p_{1}^{*})$.
@@
Under garbling for binary $Y$, we simplify notation by parameterizing $\rho$ as $\rho(0)=p_{0}$ and $\rho(1)=p_{1}$. Let $p_{0}(\varepsilon),p_{1}(\varepsilon)$ denote the values of these parameters that maximize $\Pi_{\text{garb}}$ for a given $\varepsilon$.
Appendix B Proofs from Section 4
Here we give full proofs from Section 4.
B.1 Proofs of Theorems 1 and 2
To prove Theorem 1, we first prove the following more detailed theorem.
Theorem 3.
For any cost distribution that satisfies Assumptions 1, 2, if the cost is bounded above by $\bar{C}$ with $f(\bar{C})>0$, then there exists an instantiation of the principal’s task value $b$ and a thresholding feature $X^{t}$ for which the agent prefers to reveal $X^{t}$. Specifically, the agent prefers to reveal $X^{t}$ if $b\in\left(\bar{C}+\frac{1-F(t)}{f(\bar{C})},t+\frac{F(t)}{f(t)}\right).$
Proof.
For thresholding feature $X^{t}$, we have the scaled conditional distributions:
$f_{0}(p)=\frac{f(p)}{1-F(t)}\operatorname{\mathbbm{1}}(p>t);\quad f_{1}(p)=\frac{f(p)}{F(p)}\operatorname{\mathbbm{1}}(p\leq t).$
From these, we can compute that
[…]
First, we show that for $b>\bar{C}+\frac{1-F(t)}{f(\bar{C})}$, the optimal price in the high cost setting is $p_{0}^{*}=\bar{C}$. The principal’s value is
$\Pi_{0}(p)=F_{0}(p)(v-p)=\frac{F(p)-F(t)}{1-F(t)}(v-p)\operatorname{\mathbbm{1}}(p>t).$
Concavity of $F(p)$ in Assumption 1 for $p\geq p^{*}$ implies that $\Pi_{0}(p)$ is also concave for $p\geq p^{*}$. Since the function $p+\frac{F_{0}(p)}{f_{0}(p)}$ is monotone increasing (Assumption 2), $\Pi_{0}(p)$ is either maximized for $p_{0}^{*}$ satisfying first order condition $b=p_{0}^{*}+\frac{F_{0}(p_{0}^{*})}{f_{0}(p_{0}^{*})}$ if such a $p_{0}^{*}$ exists; or $p_{0}^{*}=\bar{C}$ if $b$ is sufficiently large. Specifically, $p_{0}^{*}=\bar{C}$ if $g_{0}^{-1}(b)>\bar{C}$, where $g_{x}(p)\coloneq p+\frac{F_{x}(p)}{f_{x}(p)}$. The lower bound on $b$ follows directly: $g_{0}^{-1}(b)>\bar{C}\iff b>\bar{C}+\frac{1-F(t)}{f(\bar{C})}$.
Next, we show that for $b<t+\frac{F(t)}{f(t)}$, the price in the low cost setting is equal to price in the concealed setting: $p_{1}^{*}=p^{*}$. The principal’s value is
$\Pi_{1}(p)=F_{1}(p)(v-p)=\frac{F(p)}{F(t)}(v-p)\operatorname{\mathbbm{1}}(p\leq t).$
Concavity of $F(p)$ in Assumption 1 at $p^{*}$ implies that $\Pi_{1}(p)$ is concave near $p^{*}$ as long as $p^{*}<t$. This is sufficient for $\Pi_{1}(p)$ to be concave near its optimum $p_{1}^{*}$, since the first order condition for $\Pi_{1}(p)$ is identical to that of $\Pi(p)$. Again since $p+\frac{F_{1}(p)}{f_{1}(p)}$ is monotone increasing, $p_{1}^{*}=g_{1}^{-1}(b)$ if $g_{1}^{-1}(b)<t$; otherwise $p_{1}^{*}=t$. The upper bound on $b$ follows as $g_{1}^{-1}(b)<t\iff b<t+\frac{F(t)}{f(t)}$. Under this condition, the optimal price $p_{1}^{*}=g_{1}^{-1}(b)=g^{-1}(b)=p^{*}$.
Putting these together, when $b\in\left(\bar{C}+\frac{1-F(t)}{f(\bar{C})},t+\frac{F(t)}{f(t)}\right),$ the price in the high cost setting is $p_{0}^{*}=\bar{C}$, and the price in the low cost setting is $p_{1}^{*}=p^{*}$. The total revealed agent value is $V_{\text{rev}}(p^{*},\bar{C})>V_{\text{rev}}(p^{*},p^{*})\implies V_{\text{rev}}(p^{*},\bar{C})>V_{\text{con}}(p^{*})$.
To complete the proof, we show that there exists $t$ such that the set $\left(\bar{C}+\frac{1-F(t)}{f(\bar{C})},t+\frac{F(t)}{f(t)}\right)$ is nonempty. The set is nonempty if
$\bar{C}+\frac{1-F(t)}{f(\bar{C})}<t+\frac{F(t)}{f(t)}\iff\bar{C}-t<-\frac{1}{f(\bar{C})}+\frac{F(t)}{f(\bar{C})}+\frac{F(t)}{f(t)}\iff\bar{C}-t<\frac{1}{f(\bar{C})}\left(\left(1+\frac{f(\bar{C})}{f(t)}\right)F(t)-1\right)$ (1)
First, let $\delta_{1}>0$, and $\delta_{1}f(\bar{C})<1$, and choose $t_{1}$ sufficiently large such that $\bar{C}-t_{1}<\varepsilon$. Next, choose $\delta_{2}>0$ sufficiently small such that $\delta_{1}<\frac{1}{f(\bar{C})}(1-\delta_{2})$, and choose $t_{2}$ sufficiently large such that $\left(\left(1+\frac{f(\bar{C})}{f(t)}\right)F(t)-1\right)>1-\delta_{2}$. Finally, choose $t_{3}=\max(t_{1},t_{2})$. Then $t_{3}$ satisfies Equation (1). ∎
Intuitively, the specific bounds hold for $t$ sufficiently high relative to $b$: for a fixed $b$ not too much higher than $\bar{C}$, the bound $b<t+\frac{F(t)}{f(t)}$ holds for $t$ sufficiently high under Assumption 2. Similarly, the lower bound on $b$ is equivalent to $F(t)>1-f(\bar{C})(b-\bar{C})$, which also holds for $t$ sufficiently high. We summarize this more simply in Theorem 1 below.
Theorem 1 (Agent prefers to reveal for high thresholds).
Under Assumptions 1 and 2, if the cost is bounded above by $\bar{C}$ with $f(\bar{C})>0$, and if $b\in\left(\bar{C},\bar{C}+\frac{1}{f(\bar{C})}\right)$, then there exists a threshold $\underline{t}$ such that for all $t>\underline{t}$, the agent prefers to reveal $X^{t}$.
Proof.
This proof follows directly from Theorem 3. Fix $b\in\left(\bar{C},\bar{C}+\frac{1}{f(\bar{C})}\right)$. Choose $t_{1}$ sufficiently large that the lower bound holds: $b>\bar{C}+\frac{1-F(t_{1})}{f(\bar{C})}\iff b-\bar{C}>\frac{1-F(t_{1})}{f(\bar{C})}$. This is possible since $b>\bar{C}$ and $f(\bar{C})>0$. This lower bound will hold for any $t>t_{1}$.
Choose $t_{2}$ sufficiently large such that $b<t_{2}+\frac{F(t_{2})}{f(t_{2})}$ the upper bound holds: $b<t_{2}+\frac{F(t_{2})}{f(t_{2})}$. This is possible since $b<\max_{t\in[0,\bar{C}]}t+\frac{F(t)}{f(t)}$ (under Assumption 2). By Assumption 2, this upper bound will hold for any $t>t_{2}$. Choose $\underline{t}=\max(t_{1},t_{2})$. ∎
Lemma 6.
For $t\in\{0,\bar{C}\}$, $\Delta(t)=0$.
Proof.
When $t=0$, $X^{t}=0$ with probability 1. Therefore, $F_{0}(c)=F(c)$, $p_{0,t}^{*}=p^{*}$, and $\theta=1$, so $V_{\text{rev}}^{t}(p_{1,t}^{*},p_{0,t}^{*})=V_{\text{con}}(p^{*})$ (for any $p_{1,t}^{*}$). A similar argument holds for $t=\bar{C}$. ∎
Theorem 2 (Agent prefers to conceal for low thresholds).
Under Assumptions 1, 2, and 3, if $f(0)>0$, then $\Delta^{\prime}(0)<0$.
Proof.
For notational convenience, let the superscript $t$ represent all principal and agent values defined in Section 3.1, but with $X=X^{t}$:
$F^{t}_{0}(c)=P(C\leq c|X^{t}=0),\;\;F^{t}_{1}(c)=P(C\leq c|X^{t}=1),\text{etc.}$.
Let $p_{0}(t),p_{1}(t)$ be defined as the principal’s optimal prices under revealed $X^{t}$,
$p_{0}(t),p_{1}(t)\in\underset{p_{0},p_{1}}{\arg\max}\Pi_{\text{rev}}^{t}(p_{0},p_{1}).$
With a bit of abuse of notation, let $V_{\text{rev}}(t)\coloneq V_{\text{rev}}^{t}(p_{0}(t),p_{1}(t))$. Then $\Delta(t)=V_{\text{rev}}(t)-V_{\text{con}}(p^{*})$. Since $V_{\text{con}}(p^{*})$ does not depend on $t$, we need only show that $V_{\text{rev}}^{\prime}(0)<0$. Differentiating this, we have
$V_{\text{rev}}^{\prime}(0)$ $=\frac{\partial}{\partial t}\left(V_{1}^{t}(p_{1}(t))F(t)\right)\Big{|}_{t=0}+\frac{\partial}{\partial t}\left(V_{0}^{t}(p_{0}(t))(1-F(t))\right)\Big{|}_{t=0}$
The first term is 0:
$\frac{\partial}{\partial t}\left(V_{1}^{t}(p_{1}(t))F(t)\right)\Big{|}_{t=0}=F(0)\frac{\partial}{\partial t}V_{1}^{t}(p_{1}(t))\Big{|}_{t=0}+f(0)V_{1}^{t}(p_{1}(t))\big{|}_{t=0}=0$
Note that $\lim_{t\to 0}p_{1}(t)=0$, which implies that $V_{1}^{t}(p_{1}(t))\big{|}_{t=0}=0$. Thus,
$V_{\text{rev}}^{\prime}(0)=\frac{\partial}{\partial t}\left(V_{0}^{t}(p_{0}(t))(1-F(t))\right)\Big{|}_{t=0}=(1-F(0))\frac{\partial}{\partial t}V_{0}^{t}(p_{0}(t))\Big{|}_{t=0}-f(0)V_{0}^{t}(p_{0}(t))\Big{|}_{t=0}=\frac{\partial}{\partial t}V_{0}^{t}(p_{0}(t))\Big{|}_{t=0}-f(0)V_{\text{con}}(p^{*})$ (2)
We solve for the above using the chain rule, substituting $p=p_{0}(t)$:
$\frac{\partial}{\partial t}V_{0}^{t}(p_{0}(t))\Big{|}_{t=0}=\frac{\partial}{\partial p}V_{0}^{t}(p)\Big{|}_{t=0}p_{0}^{\prime}(0)+\frac{\partial}{\partial t}V_{0}^{t}(p)\Big{|}_{t=0}$ (3)
Solving for each of the terms in this expression:
$\frac{\partial}{\partial p}V_{0}^{t}(p)\Big{|}_{t=0}=F_{0}^{t}(p)\big{|}_{t=0}=F(p^{*})$ (4)
[…]
$\frac{\partial}{\partial t}F_{0}^{t}(p)\Big{|}_{t=0}$ $=\frac{\partial}{\partial t}\frac{F(p)-F(t)}{1-F(t)}\Big{|}_{t=0}=\frac{f(t)(F(p)-1)}{(1-F(t))^{2}}\Big{|}_{t=0}=f(0)(F(p^{*})-1)$
[…]
$\implies\frac{\partial}{\partial t}V_{0}^{t}(p)\Big{|}_{t=0}=f(0)\left(p^{*}(F(p^{*})-1)-E[\operatorname{\mathbbm{1}}(C<p^{*})C]\right)=f(0)\left(V(p^{*})-p^{*}\right)$ (5)
Finally, to solve for $p_{0}^{\prime}(t)$, we either have $p_{0}^{\prime}(t)=0$ for $b>>C$, in which case the proof is complete and it immediately follows that $\Delta^{\prime}(0)<0$; or, we suppose $p_{0}(t)$ satisfies the first order condition: $p_{0}(t)+\frac{F(p_{0}(t))-F(t)}{f(p_{0}(t))}=v$. Differentiating on both sides of this first order condition, we have
$p_{0}^{\prime}(t)=\frac{f(t)f(p_{0}(t))}{2f(p_{0}(t))^{2}-F(p_{0}(t))f^{\prime}(p_{0}(t))+F(t)f^{\prime}(p_{0}(t))}\implies p_{0}^{\prime}(0)=\frac{f(0)}{2f(p^{*})-\frac{F(p^{*})}{f(p^{*})}f^{\prime}(p^{*})}$ (6)
We now combine Equations (4), (5), and (6) with Equation (3) to get that
$\frac{\partial}{\partial t}V_{0}^{t}(p_{0}(t))\Big{|}_{t=0}=f(0)\left(\frac{F(p^{*})}{2f(p^{*})-\frac{F(p^{*})f^{\prime}(p^{*})}{f(p^{*})}}-\left(V_{\text{con}}(p^{*})-p^{*}\right)\right).$
Putting this all together with Equation (2), $V_{\text{rev}}^{\prime}(0)<0$ if and only if
$f(0)\left(\frac{F(p^{*})}{2f(p^{*})-\frac{F(p^{*})f^{\prime}(p^{*})}{f(p^{*})}}-\left(V_{\text{con}}(p^{*})-p^{*}\right)-V_{\text{con}}(p^{*})\right)<0$
$\iff 1-\frac{F(p^{*})}{p^{*}f(p^{*})}+\frac{\partial}{\partial p}\frac{F(p)}{f(p)}\Big{|}_{p=p^{*}}>0$
This holds by Assumptions 2 and 3. Therefore, $\Delta^{\prime}(0)<0$.
B.2 Proofs from Section 4.2
Lemma 1.
Total welfare increases under revelation ($W_{\text{rev}}(\rho^{*})-W_{\text{con}}(p^{*})$) only if task completion quantity increases under revelation,
$F(p^{*})<\mathbb{E}[\operatorname{\mathbbm{1}}(C<\rho^{*}(X)].$
Proof.
Let $\underline{\mathcal{X}}\subseteq\mathcal{X}$ be the set of $x$ values for which $\rho(x)<p^{*}$, and let $\overline{{\mathcal{X}}}={\mathcal{X}}\setminus\underline{\mathcal{X}}$.
$W_{\text{rev}}(\rho^{*})-W_{\text{con}}(p^{*})$ $=\mathbb{E}[\operatorname{\mathbbm{1}}(C<\rho^{*}(X))(b-C)]-\mathbb{E}[\operatorname{\mathbbm{1}}(C<p^{*})(b-C)]$
$=\mathbb{E}[\operatorname{\mathbbm{1}}(C<\rho^{*}(X))(b-C)|X\in\underline{\mathcal{X}}]{\mathbb{P}}(X\in\underline{\mathcal{X}})+\mathbb{E}[\operatorname{\mathbbm{1}}(C<\rho^{*}(X))(b-C)|X\in\overline{{\mathcal{X}}}]{\mathbb{P}}(X\in\overline{{\mathcal{X}}})-\mathbb{E}[\operatorname{\mathbbm{1}}(C<p^{*})(b-C)]$
$=\mathbb{E}[(\operatorname{\mathbbm{1}}(C<\rho^{*}(X))-\operatorname{\mathbbm{1}}(C<p^{*}))(b-C)|X\in\overline{{\mathcal{X}}}]{\mathbb{P}}(X\in\overline{{\mathcal{X}}})$
$\quad-\mathbb{E}[(\operatorname{\mathbbm{1}}(C<p^{*})-\operatorname{\mathbbm{1}}(C<\rho^{*}(X)))(b-C)|X\in\underline{\mathcal{X}}]{\mathbb{P}}(X\in\underline{\mathcal{X}})$
$\leq\mathbb{E}[(\operatorname{\mathbbm{1}}(C<\rho^{*}(X))-\operatorname{\mathbbm{1}}(C<p^{*}))(b-p^{*})|X\in\overline{{\mathcal{X}}}]{\mathbb{P}}(X\in\overline{{\mathcal{X}}})$
$\quad-\mathbb{E}[(\operatorname{\mathbbm{1}}(C<p^{*})-\operatorname{\mathbbm{1}}(C<\rho^{*}(X)))(b-p^{*})|X\in\underline{\mathcal{X}}]{\mathbb{P}}(X\in\underline{\mathcal{X}})$
$\implies W_{\text{rev}}(\rho^{*})-W_{\text{con}}(p^{*})\leq(b-p^{*})(\mathbb{E}[\operatorname{\mathbbm{1}}(C<\rho^{*}(X)]-\mathbb{E}[\operatorname{\mathbbm{1}}(C<p^{*})])$
Since $(b-p^{*})>0$, if $(\mathbb{E}[\operatorname{\mathbbm{1}}(C<\rho^{*}(X)]-\mathbb{E}[\operatorname{\mathbbm{1}}(C<p^{*})])\leq 0$, then $W_{\text{rev}}(\rho^{*})-W_{\text{con}}(p^{*})\leq 0$. ∎
Appendix C Proofs from Section 5
C.1 Proofs from Section 5.4
Lemma 5 (Privacy of garbling mechanism).
For $\theta\in(0,1)$, for all $\varepsilon\leq O(\theta)$, the garbling mechanism is an $O(\varepsilon)$-local randomizer with respect to $X$.
Proof.
Suppose without loss of generality that $\theta<\frac{1}{2}$ (i.e., $X=0$ with higher probability).
$\max_{y,x\in\{0,1\}}\frac{{\mathbb{P}}(Y=y|X=x)}{{\mathbb{P}}(Y=y|X=1-x)}\leq\frac{{\mathbb{P}}(Y=1|X=1)}{{\mathbb{P}}(Y=1|X=0)}=\frac{\theta(1-\varepsilon)+\varepsilon}{\theta(1-\varepsilon)}\leq e^{O(\varepsilon)}$
The last inequality holds for $\varepsilon$ sufficiently small: $\varepsilon\leq\frac{\theta}{\theta+1}\implies\frac{\varepsilon}{\theta(1-\varepsilon)}\leq 1$.
A pricing mechanism that accesses $X$ only through this garbling mechanism is therefore locally differentially private with a corresponding privacy parameter. ∎
C.2 Proofs from Section 5.3
Lemma 3 (Reducing garbling initially increases total welfare).
Proof.
Lemma 2 (Optimal garbling increases welfare over concealment).
Proof.
Proof.
Appendix D Additional Results on Welfare Effects of Information Revelation
D.1 Uniform Simulation for Thresholding Features
To illustrate Theorems 1 and 2 on thresholding features, let $C\sim\text{Unif}(0,1)$. Let $X^{t}$ be a thresholding feature for $C$, where $X^{t}=1$ if $C\leq t$, and 00 otherwise. Figure 5 shows the agent’s utility difference between revealing and concealing $X^{t}$ for all pairs of $b\in(0,2),t\in(0,1)$.
{tblr}
p0.1cmc &$V_{\text{rev}}(p_{0}^{*},p_{1}^{*})-V_{\text{con}}(p^{*})$
[…]
$t$
Figure 5: Difference between agent’s utility in the revealed setting and concealed settings when $C\sim\text{Unif}(0,1)$, and $X=X^{t}$ is a thresholding feature. For a given pair $(b,t)$, a positive utility difference means the agent prefers to reveal $X^{t}$, and a negative utility difference means the agent prefers to conceal $X^{t}$. The agent always prefers to conceal for $t$ sufficiently close to 00, which matches Theorem 2. For $b\in(1,2)$, the agent prefers to reveal for $t>\frac{b}{2}$, which matches Theorem 1. At $b=2$, the agent always prefers to conceal. For $b<1$, the agent actually prefers to conceal for sufficiently high $t$.
D.2 Agent’s Revelation Incentives: General Conditions
We give more general conditions the agent to prefer to reveal or conceal for $X$ beyond thresholding features.
We first show that the agent prefers the hidden information setting if the cost conditional on $X$ is close to zero for some values of $X$, and still not very high for other values of $X$. Theorem 1 shows a general hazard rate condition that leads to this conclusion.
Next, we give sufficient conditions for the agent to conceal and reveal under general $X$, applying identities related to the analysis done by Aguirre et al. [2010].
D.2.1 Concealment condition with one zero-cost type
@@
We present a sufficient condition for the agent to prefer to conceal $X$ when one of the agent types is anchored at zero. That is, $X=1$ implies that the agent incurs zero cost. Proposition 1 gives a sufficient condition on $F_{0}$ for the agent to prefer for the environmental variable $X$ to remain concealed.
@@
Intuitively, the elasticity $\eta$ captures the sensitivity of task completion to the offered price. Thus, the inequality in equation (7) corresponds to a scenario when the sensitivity of the task completion to price when $X=0$ does not differ too strongly from that of the concealed setting. For example, this arises when $F_{0}$ is close to the constant function 1. We give another example using the exponential distribution in Section D.3 below, where equation (7) holds if the mean of $F_{0}$ is low enough. In summary, the agent prefers concealment if the higher cost type still has relatively low cost.
D.2.2 Concealment and revelation conditions under a decreasing ratio assumption
To give additional sufficient conditions for concealment and revelation for variables more general than thresholding features, we apply an analysis technique similar to that of Aguirre et al. [2010], who analyzed the effects of third degree monopoly price discrimination on total welfare.